A 8 Ohms loudspeaker and a 8 Ohms loudspeaker are connected in parallel across the terminals of an amplifier. Assuming the speakers behave as resistors, what is the equivalent resistance of the two speakers?

_____________Ohms
1
Expert's answer
2012-02-14T09:55:59-0500
Because 8 Ohms loudspeaker and a 8 Ohms& loudspeaker are connected in parallel across the terminals of an amplifier and they are behaving as resistors we can apply formula for resistances in case of parallel connection:
&
1/R = 1/R1 + 1/R2 = 1/8+1/8 = 1/4 ==> R = 4 Ohms. - answer